# Dream to Fly: Model-Based Reinforcement Learning for Vision-Based Drone Flight

### v1

@@ Abstract
Autonomous drone racing has risen as a challenging robotic benchmark for testing the limits of learning, perception, planning, and control. Expert human pilots are able to agilely fly a drone through a race track by mapping the real-time feed from a single onboard camera directly to control commands. Recent works in autonomous drone racing attempting direct pixel-to-commands control policies (without explicit state estimation) have relied on either intermediate representations that simplify the observation space or performed extensive bootstrapping using Imitation Learning (IL). This paper introduces an approach that learns policies from scratch, allowing a quadrotor to autonomously navigate a race track by directly mapping raw onboard camera pixels to control commands, just as human pilots do. By leveraging model-based reinforcement learning (RL)—specifically DreamerV3—we train visuomotor policies capable of agile flight through a race track using only raw pixel observations. While model-free RL methods such as PPO struggle to learn under these conditions, DreamerV3 efficiently acquires complex visuomotor behaviors. Moreover, because our policies learn directly from pixel inputs, the perception-aware reward term employed in previous RL approaches to guide the training process is no longer needed. Our experiments demonstrate in both simulation and real-world flight how the proposed approach can be deployed on agile quadrotors. This approach advances the frontier of vision-based autonomous flight and shows that model-based RL is a promising direction for real-world robotics.
@@ I INTRODUCTION
In recent years, quadrotors have become a central focus of robotics research, emerging as versatile platforms with untapped potential across multiple domains, including search and rescue, inspection, agriculture, cinematography, delivery, passenger air vehicles, and space exploration. Among these applications, the field of autonomous drone racing has attracted particular interest in the research community [1], showcasing the remarkable progress in agile flight control. This domain not only benefits from cutting-edge robotics research but also pushes the limits of what is possible [2] by challenging these flying machines to outperform the most skilled human pilots, as shown by recent successes against the world’s best pilots [3, 4]. The ability to navigate complex race tracks at high speeds, making split-second decisions, not only pushes the boundaries of what is possible in autonomous flight, but also stresses the capabilities of machine learning, perception algorithms, planning and control systems, requiring them to perform at peak efficiency under highly dynamic and demanding conditions.
@@
Traditionally, most autonomous drone racing systems have relied on explicit state estimation integrating data from inertial measurement units (IMUs) and other onboard sensors to maintain stability and optimize performance [5, 6, 7, 8, 9, 3, 10, 11].
@@
Figure 1: Real-world deployment of our DreamerV3 policy in the Figure 8 track. During training, the agent learns a world model from interactions with the environment. At the same time, the actor-critic policy is trained by sampling the predictions of the world model, also called imagination. The onboard images consumed by the network are shown in red.
@@
However, professional human pilots rely solely on visual feedback from an onboard camera, showcasing a remarkable ability to navigate complex environments only from visual input extracted from an unique onboard camera. Emulating this human ability to fly based solely on visual information remains a significant challenge for autonomous systems.
@@
Bridging the gap between perception and control – learning directly from pixels to actions without the need for explicit state estimation – remains largely unfulfilled. While reinforcement learning (RL) has shown promise in various robotic applications, applying it to vision-based tasks introduces unique difficulties, especially in real-time environments such as drone racing.
@@
One of the most salient and recent works in this domain [12] successfully manages to fly a drone through a race track using only visual inputs as observations. In this case, the visual inputs are first preprocessed and distilled into a simpler intermediate representation in the form of a binary mask where only the racing gates are visible. This intermediate representations reduces the overall observation space complexity and allows the RL policy to learn the behavior efficiently. However, this work relies on a simplified observation space: the raw visual input is preprocessed into a binary mask highlighting only the gates. This simplification, while reducing the complexity of the observation space and facilitating learning, also discards valuable information. For instance, excluding background information can hinder the agent’s ability to navigate when no gates are immediately visible, or obtaining the gravity direction from the horizon line, for example. These challenges to learn directly from pixels primarily arise due to the high dimensionality of the observation space and the sample inefficiency of model-free RL methods, which often require extensive interaction with the environment to learn effectively. Another recent work in flying a quadrotor only from pixels is  [13], but in this case extensive bootstrapping with Imitation Learning is performed.
@@
Model-based reinforcement learning (MBRL) presents a promising solution to these challenges by learning the transition dynamics of the system, called the world model, and using this it to predict and optimize future actions. MBRL is generally more sample efficient than its model-free counterpart, reducing the need for extensive environment interaction. It also generalizes well to different tasks as the transition model does not vary vastly across different tasks in the same system description. However, this approach typically comes at the cost of increased complexity and longer training times, as the system must simultaneously learn a world model and optimize a control policy. Additionally, the inherent complexity of model-based RL architectures, with their multiple interdependent components, often makes them more difficult to tune and optimize.
@@
Recent works in MBRL, such as the DreamerV3 architecture [14], have helped mitigate these issues, offering a model-based RL approach that is more accessible in terms of tuning and optimization. Despite its successes in controlled environments, however, the DreamerV3 framework has so far seen limited application in real-world robotic systems, especially when it comes to real-world robotics tasks involving direct pixel-to-action learning.
This paper aims to address these gaps by presenting an approach for vision-based autonomous quadrotor flight, utilizing model-based reinforcement learning to directly map pixel observations to control commands. We train a DreamerV3 architecture to perform the task of drone racing with the same observations and control input modality as human pilots: raw pixels and collective thrust and bodyrates, respectively. We show that while PPO (Proximal Policy Optimization [15]) cannot learn directly from pixels to commands, our DreamerV3 approach efficiently learns to fly agilely through gates, in both simulation and in the real-world. This is the first approach that is able to learn to race directly from pixels, without intermediate representation and without explicit state estimation.
Contributions
This paper introduces the following contributions:
  • To the best of our knowledge, we present the first autonomous agent that is able to fly a quadrotor platform uniquely using pixels to commands, without intermediate representations and without explicit state estimation
  • We show that model-free RL, in particular PPO, cannot learn this behavior in an end-to-end fashion.
@@
  • We show that our agents exhibit emergent behavior that naturally guides the drone camera view towards feature rich areas, such as the next gates, without requiring handcrafted shaped reward terms to incentivize gate-viewing as done in previous work.
  • We deploy our system in both simulation and real-world, further validating the applicability of MBRL to real-world mobile robots.
@@ II-A Reinforcement Learning from Pixels
In recent years, RL algorithms have achieved incredible feats in simple environments learning directly from pixels, such as arcade games, used as sample efficient simulators to benchmark the capabilities of AI agents [16]. This success in RL has been repeatedly shown in environments in growing complexity, such as beating human performance in Atari games  [17] or in more complex games such as Go [18]. However, these successes have been primarily in environments where the action space is discrete, and not in continous action spaces. This gap is addressed by the DeepMind Control Suite [19], which presents a diverse set of environments with different observation spaces and different action modalities — both continuous and discrete. However, learning directly from pixels in these scenarios has been shown to be less sample efficient, yielding suboptimal performance when compared to state-based learning, even in simulation environments. Recent efforts have tried to improve this by building on top of intermediate visual representations [20, 21, 22, 23, 12]. For example, previous work explored different kinds of state representations via various supervised and self-supervised learning methods, such as auto-encoders [20], future predictions [22], contrastive unsupervised representations [23] or semantic segmentation [12], with the objective of reducing the performance difference between state-based and pixel-based RL.
@@
In order to address the challenge of learning complex behaviours efficiently directly from pixels, some methods have focused on data augmentation [24, 25]. Policies that map camera observations directly to commands — also called sensorimotor or visuomotor policies — are mostly learned through extensive data simulation, making use of domain randomization techniques or incorporating additional priviledge information such as joint angles  [26, 27, 28] For more complicated tasks, the robotics community has resorted to the usage of expert demonstrations, marking a shift towards imitation learning approaches, contrary to the learning from scratch using RL paradigm. This is showcased by recent works, most notably  [29].
II-B Vision-based flight
In the last decade, a plethora of works that aim to fly a drone from pixels to high level commands has emerged [6, 30, 31, 32, 33]. These approaches are not using reinforcement learning, and try to tackle the navigation task in the more general sense, therefore far from agile flight. In [34] the authors map features and IMU to control commands to fly aggressive maneuvers, learning via imitation learning. In the realm of state-based policy learning, using RL for the low-level control of quadrotors has been demonstrated to perform better than classical methods  [35, 36]. Particularly, in state-based drone racing, RL algorithms have been able to demonstrate agile flight [37, 38], surpassing even state-of-the-art optimal control methods [4]. The adaptability and versatility of RL methods has been showing more and more success in vision-based flight in several recent works. One of the earliest uses of RL for vision-based quadrotor navigation is [39], where a CAD model was used to train an RL policy that would output discrete velocity commands (’forward’, ’right’, ’left’). This was enabled thanks to the onboard stabilizing controller and explicit state estimation through a VIO pipeline.
More recently and for the first time, a quadrotor that combines VIO with an RL controller has been able to win a fair drone race against the top human world champions [3]. Even though this represents a remarkable achievement and a landmark for robotics, the drone is still a custom built research drone [40], which runs a state-estimation pipeline on board. On top of this, the policies needed to be fine-tuned by using real-world data captured with a motion-capture system. Therefore, the observation modality is still not the same that the humans have, which is exclusively a single monocular feed from the onboard camera of the platform. In a more recent work [12], the authors were able for the first time to achieve agile flight without explicit state estimation. The authors convert the pixel image into a black and white image that contains only the gates, and from this intermediate representation they train a model-free RL policy from scratch. Even though in that case they were able to fly through the gates, they still were not able to train directly from the camera feed, and the policy is not aware of any of the texture or background information.
In this work, we aim to overcome this limitation by using model-based RL (DreamerV3 [14]) to train a policy that can map directly RGB pixels to control commands, just like humans do.
II-C Model-based Reinforcement Learning for Real-World Robots
@@
As mentioned above, there are many approaches that have made significant progress in learning policies directly from pixels. However, the landscape of model-based RL (MBRL) methods applied to real-world robotics remains relatively limited, and existing research primarily focuses on simulation environments. Within the context of real-world robotic applications, prior MBRL work often relies on state estimations rather than raw pixel observations. Specifically, there are some works tackling the problem of flying a quadcopter using model-based RL, for high-level control [41] or low-level controller learning [36]. However, these works have been using state estimation as observations, and not directly pixels as observations.
@@
When speaking about works that use MBRL from pixels and that deploy in real-world robots, the number of works is even lower. One of the early works in this direction is [42], where a learned world model of the dynamics is built from pixels, and is then used by an MPC planner to get the optimal policy, which outputs torque commands for a real-world manipulator. Another interesting work that applies DreamerV3 to a real-world robotic system is [43], where the table top labyrinth game is solved in the real world by using top-view fixed camera in conjunction with the measurement of the table angle. However, this work does not belong to the mobile robotics category. One of the most salient works that applies DreamerV3 [14] to real-world robotics is [44], where different policies are trained from images, depth and state, for different tasks. These include pick-and-place manipulation utilizing a fusion of RGB images, depth data, and proprioceptive sensing, as well as quadruped locomotion trained solely on proprioceptive sensor data. Notably, the only instance of direct learning from pixels to commands involves a 2D navigation task with a Sphero robot, a relatively simple scenario in terms of dynamics and environmental complexity.
@@
Figure 2: The process begins with data collection in the real environment using the current policy, storing experiences in a replay buffer. This buffer is used to train the world model components: the encoder, decoder, RSSM, dynamics model, reward model, and continue predictor (Section III-E). Subsequently, an actor-critic policy is trained within the learned world model to maximize expected (imagined) returns. This updated policy is then used to collect new data, restarting the loop.
@@ III-A Problem Statement
Our objective is to navigate a quadrotor through a series of gates in minimum time while avoiding collisions with the environment and the gates. Formally, we seek a policy $\pi(x)$ that maps raw visual observations directly to control commands, minimizing the time required to complete the course. We address this challenge using reinforcement learning (RL), framing the problem within the Markov Decision Process (MDP) framework.
III-B MDP formulation
The Markov Decision Process (MDP) formalism provides a structured approach for modeling decision-making problems. An MDP is defined by the tuple $(\mathcal{X},\mathcal{A},P,R,\gamma)$, where $\mathcal{X}$ is the set of possible states, $\mathcal{A}$ is the set of possible actions, $P(x_{k+1}|x_{k},a_{k})$ is the state transition probability, $R(x_{k},a_{k})$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. In our vision-based drone racing task, the state $x_{k}$ is represented by the raw image captured by the onboard camera at time step $k$. The action $a_{k}$ consists of the collective thrust and body rates applied to the quadrotor. The reward function is designed to incentivize fast and collision-free navigation through the gates (detailed in Section III-D) The objective in an MDP is to find an optimal policy $\pi_{\theta}^{*}:\mathcal{X}\rightarrow\mathcal{A}$ that maximizes the expected cumulative discounted reward by optimizing the parameters $\theta$ of a neural network,
$\pi_{\theta}^{*}=\arg\max_{\pi}\mathbf{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r(k)\right].$
@@ III-B MDP formulation
III-C Observation and Action Spaces
@@ III-D Reward Function
$r(k)=\begin{cases}-4.0,&\text{if collision;}\\ +10,&\text{if gate passed;}\\ \begin{aligned} b_{1}(&\|\mathbf{g}_{k}-\mathbf{p}_{k-1}\|-\|\mathbf{g}_{k}-\mathbf{p}_{k}\|)\\ &-b_{2}\|\bm{\omega}_{k}\|,\end{aligned}&\text{otherwise,}\end{cases}$ (1)
where $\mathbf{g}_{k}$ is the center of the target gate, $\mathbf{p}_{k}$ and $\mathbf{p}_{k-1}$ are the drone’s positions at the current and previous time steps, respectively, and $\bm{\omega}_{k}$ is the body rate vector. The primary component of the reward, $b_{1}(\|\mathbf{g}_{k}-\mathbf{p}_{k-1}\|-\|\mathbf{g}_{k}-\mathbf{p}_{k}\|)$, is called the progress term, as it encourages the drone to move closer to the target gate. The term $b_{2}\|\bm{\omega}_{k}\|$ penalizes excessive body rates. We use the coefficients $b_{1}=1.0,b_{2}=0.01$. This formulation directly rewards progress toward the gate center. Importantly, deviations from the exact path defined by the waypoints are not penalized, allowing the agent to discover potentially more efficient trajectories.
In addition to the continuous reward component, discrete rewards are provided for specific events. A large negative reward $(-4)$ is given upon collision with the environment or a gate, terminating the episode. A positive reward $(+10)$ is awarded when the drone successfully passes through a gate, triggering the transition to the next target gate. An additional positive reward $(+10)$ is given upon finishing the race.
III-E Model-based Reinforcement Learning: DreamerV3

### v2

@@ Abstract
Autonomous drone racing has risen as a challenging robotic benchmark for testing the limits of learning, perception, planning, and control. Expert human pilots are able to fly a drone through a race track by mapping pixels from a single camera directly to control commands. Recent works in autonomous drone racing attempting direct pixel-to-commands control policies have relied on either intermediate representations that simplify the observation space or performed extensive bootstrapping using Imitation Learning (IL). This paper leverages DreamerV3 to train visuomotor policies capable of agile flight through a racetrack using only pixels as observations. In contrast to model-free methods like PPO or SAC, which are sample-inefficient and struggle in this setting, our approach acquires drone racing skills from pixels. Notably, a perception-aware behaviour of actively steering the camera toward texture-rich gate regions emerges without the need of handcrafted reward terms for the viewing direction. Our experiments show in both, simulation and real-world flight using a hardware-in-the-loop setup with rendered image observations, how the proposed approach can be deployed on real quadrotors at speeds of up to 9 m/s. These results advance the state of pixel-based autonomous flight and demonstrate that MBRL offers a promising path for real-world robotics research.
@@ I INTRODUCTION
In recent years, quadrotors have become a central focus of robotics research, emerging as versatile platforms with untapped potential across multiple domains, including search and rescue, inspection, agriculture, cinematography, delivery, passenger air vehicles, space exploration [22] and drone racing [13]. The drone racing domain not only benefits from cutting-edge robotics research [23] but also pushes the limits of what is possible by challenging these flying machines to outperform the most skilled human pilots, as shown by recent successes against the world’s best pilots [15, 33].
@@
Traditionally, most autonomous drone racing systems have relied on explicit state estimation integrating data from inertial measurement units (IMUs) and other onboard sensors to maintain stability and optimize performance [14, 16, 5, 26, 15].
@@
Figure 1: Real-world deployment of our DreamerV3 policy in the Figure 8 track. During training, the agent learns a world model from interactions with the environment. At the same time, the actor-critic policy is trained by sampling the predictions of the world model, also called imagination. The rendered images consumed by the network are marked in red.
@@
However, professional human pilots rely solely on visual feedback from a single onboard camera, showcasing a remarkable ability to navigate complex environments purely from visual inputs. Emulating this human ability to fly based solely on visual information remains a significant challenge for autonomous systems.
@@
Closing the loop between perception and control – learning directly from pixels to actions, without the need for explicit state estimation – remains largely unfulfilled. While reinforcement learning (RL) has shown promise in various robotic applications, applying it to vision-based tasks introduces unique difficulties in robotics.
@@
One of the most recent works in this domain [8] manages to train a Proximal Policy Optimization (PPO) [30] policy to fly a drone through a race track from binary image representations. In this case, the visual inputs are first preprocessed and distilled into a simpler intermediate representation in the form of a binary mask where only the racing gates are visible. This intermediate representation reduces the overall observation space complexity and allows the PPO policy to learn the behavior efficiently. However, this work relies on a simplified observation space: the raw visual input is preprocessed into a binary mask highlighting only the gates. This simplification, while reducing the complexity of the observation space and facilitating learning, also discards valuable information. For instance, excluding background information can hinder the agent’s ability to navigate when no gates are immediately visible, or to obtain the gravity direction from the horizon line, for example. In [39] the authors present another recent work that tackles quadrotor flight from pixel observations. However, it relies on IL bootstrapped from a pre-trained expert policy, which had access to the full state and introduces a strong dependency on privileged information. Despite their success, one major challenge in robot learning is the need for large amounts of physical interaction data, which can be very expensive and challenging to obtain. This challenge is compounded in vision-based control tasks, where the high dimensionality of image observations further increases data requirements.
@@
MBRL presents a promising solution to these challenges because it is generally more sample efficient than its model-free counterpart, reducing the need for extensive environment interactions. MBRL learns the transition dynamics of the system, known as world model, and uses it to predict and optimize future actions. It also generalizes well to different tasks as the transition model does not vary vastly across different tasks in the same system description. However, despite the sample efficiency, this approach typically comes at the cost of longer training times, as the system must simultaneously learn a world model and optimize a control policy. Additionally, the inherent complexity of model-based RL architectures, with their multiple interdependent components, often makes them more difficult to tune and optimize.
@@
Recent works in MBRL, such as the DreamerV3 architecture [12], have helped mitigate these issues, offering an approach that is more accessible in terms of tuning and optimization. Although DreamerV3 has shown strong results in controlled settings, its adoption in real-world robotic systems, particularly for pixel-based tasks, remains limited. We build on DreamerV3 to demonstrate its effectiveness in vision-based robotics tasks that require egocentric perception and agile control, such as drone racing. Our contributions are as follows:
  • Pixel-to-command MBRL system for quadrotor flight: We train quadrotor policies from scratch, without relying on intermediate representations or bootstrapping from imitation learning, which were necessary in previous works.
@@
  • Emergent perception-aware behaviour: the drone’s camera view is naturally guided towards feature-rich areas such as the next gates. This behavior arises directly from our end-to-end optimization from pixels to commands, without requiring handcrafted reward terms that were common in previous methods.
  • Real-world policy transfer: We validate our learned policies by deploying them on a physical quadrotor with a hardware-in-the-loop (HIL) setup.
Our approach demonstrates that the learned policy effectively controls real-world dynamics directly from rendered pixel observations, and further validates the applicability and potential of MBRL for real-world mobile robotic tasks.
@@ II-A Reinforcement Learning from Pixels
In recent years, RL algorithms have achieved incredible feats in simple environments learning directly from pixels, such as arcade games, used as sample efficient simulators to benchmark the capabilities of AI agents [3]. This success in RL has been repeatedly shown in environments in growing complexity, such as beating human performance in Atari games  [21] or in more complex games such as Go [31]. However, these successes have been primarily in environments where the action space is discrete, and not in continuous action spaces. This gap is addressed by the DeepMind Control Suite [35], which presents a diverse set of environments with different observation spaces and different action modalities — both continuous and discrete. However, learning directly from pixels in these scenarios has been shown to be less sample efficient, yielding suboptimal performance when compared to state-based learning, even in simulation environments. Recent efforts have tried to improve this by building on top of intermediate visual representations [41, 1, 10, 19, 8]. For example, previous work explored different kinds of state representations via various supervised and self-supervised learning methods, such as auto-encoders [41], future predictions [10], contrastive unsupervised representations [19] or semantic segmentation [8], with the objective of reducing the performance difference between state-based and pixel-based RL.
@@
In order to address the challenge of learning complex behaviors efficiently directly from pixels, some methods have focused on data augmentation [18, 40]. Policies that map camera observations directly to commands — also called sensorimotor or visuomotor policies — are mostly learned through extensive data simulation, making use of domain randomization techniques or incorporating additional privileged information such as joint angles  [28, 36, 20] For more complicated tasks, the robotics community has resorted to the usage of expert demonstrations, marking a shift towards imitation learning approaches, contrary to the learning from scratch using RL paradigm. This is showcased by recent work, most notably  [7].
II-B Model-based RL for Real-World Robotics
@@
As mentioned above, there are many approaches that have attempted to train policies using pixels as observations, relying in simplifying assumptions. However, the landscape of model-based RL (MBRL) methods applied to real-world robotics remains relatively limited, and existing research primarily focuses on simulation environments. Within the context of real-world robotic applications, prior MBRL work often relies on state estimations rather than raw pixel observations. Specifically, there are some works tackling the problem of flying a quadcopter using model-based RL, for high-level control [2] or low-level controller learning [17]. However, these works have been using state estimation as observations, and not directly pixels as observations.
@@
When speaking about works that use MBRL from pixels and that deploy in real-world robots, the number of works is even lower. One of the early works in this direction is [37], where a learned world model of the dynamics is built from pixels, and is then used by an MPC planner to get the optimal policy, which outputs torque commands for a real-world manipulator. Another interesting work that applies DreamerV3 to a real-world robotic system is [4], where the table top labyrinth game is solved in the real world by using top-view fixed camera in conjunction with the measurement of the table angle. However, this setup does not fall under mobile robotics, as the observations are not egocentric but rather taken from a fixed viewpoint with most of the environment remaining unchanged in the field of view. One of the most salient works that applies DreamerV3 [12] to real-world robotics is [38], where different policies are trained from images, depth and state, for different tasks. These include pick-and-place manipulation utilizing a fusion of RGB images, depth data, and proprioceptive sensing, as well as quadruped locomotion trained solely on proprioceptive sensor data. Notably, the only instance of direct learning from pixels to commands involves a 2D navigation task with a Sphero robot, a relatively simple scenario in terms of dynamics and environmental complexity.
@@
Figure 2: The process begins with data collection in the simulation environment using the current policy, storing experiences in a replay buffer. This buffer is used to train the world model components: the encoder, decoder, RSSM, dynamics model, reward model, and continue predictor (Section III-D). Subsequently, an actor-critic policy is trained within the learned world model to maximize expected (imagined) returns. This updated policy is then used to collect new data, restarting the loop.
@@ III-A Problem Statement
Formally, we seek a policy $\pi(x)$ that maps raw visual observations $x$ directly to control commands $a$, minimizing the time required to complete the course. We address this challenge using reinforcement learning (RL), framing the problem within the Partially Observable Markov Decision Process (POMDP) framework. A POMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{X},P,R,\gamma,\Omega)$, where $\mathcal{S}$ is the set of latent states, $\mathcal{A}$ is the set of actions, $\mathcal{X}$ is the set of observations, $P(s_{k+1}\mid s_{k},a_{k})$ is the state transition probability, $R(s_{k},a_{k})$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\Omega(x_{k}\mid s_{k})$ is the observation model. In our vision-based drone racing task, the latent state $s_{k}$ is not observed directly; instead, the agent receives the RGB image observation $x_{k}\in\mathcal{X}$ rendered at time step $k$. The action $a_{k}$ consists of the collective thrust and body rates applied to the quadrotor. The reward function is designed to incentivize fast and collision-free navigation through the gates (detailed in Section III-C) The objective in a POMDP is to find an optimal policy $\pi_{\theta}^{*}:\mathcal{X}\rightarrow\mathcal{A}$ that maximizes the expected cumulative discounted reward by optimizing the parameters $\theta$ of a neural network,
$\pi_{\theta}^{*}=\arg\max_{\pi}\mathbf{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}\right].$
@@
III-B Observation and Action Spaces
@@ III-C Reward Function
$r_{k}=\begin{cases}r_{collision},&\text{if collision;}\\ r_{passed},&\text{if gate passed;}\\ \begin{aligned} b_{1}(&\|\mathbf{g}_{k}-\mathbf{p}_{k-1}\|-\|\mathbf{g}_{k}-\mathbf{p}_{k}\|)\\ &-b_{2}\|\bm{\omega}_{k}\|,\end{aligned}&\text{otherwise,}\end{cases}$ (1)
where $r_{k}$ is the reward at timestep $k$, $\mathbf{g}_{k}$ is the center of the target gate, $\mathbf{p}_{k}$ and $\mathbf{p}_{k-1}$ are the drone’s positions at the current and previous time steps, respectively, and $\bm{\omega}_{k}$ is the body rate vector. The primary component of the reward, $b_{1}(\|\mathbf{g}_{k}-\mathbf{p}_{k-1}\|-\|\mathbf{g}_{k}-\mathbf{p}_{k}\|)$, is called the progress term, as it encourages the drone to move closer to the target gate. The term $b_{2}\|\bm{\omega}_{k}\|$ penalizes excessive body rates. We use the coefficients $b_{1}=1.0,b_{2}=0.01$. This formulation directly rewards progress toward the gate center. Importantly, deviations from the exact path defined by the waypoints are not penalized, allowing the agent to discover potentially more efficient trajectories.
In addition to the continuous reward component, discrete rewards are provided for specific events, e.g., $r_{collision}=-4.0$ and $r_{passed}=+10.0$.
III-D Model-based Reinforcement Learning: DreamerV3